1. Break-Even Point

Radison Inc. sells a product for $97 per unit. The variable cost is $52 per unit, while fixed costs are $516,375.

Determine (a) the break-even point in sales units and (b) the break-even point if the selling price were increased to $103 per unit.

a. Break-even point in sales units    ? units
b. Break-even point if the selling price were increased to $103 per unit ? units

2. Outdoors Company sells a product for $150 per unit. The variable cost is $65 per unit, and fixed costs are $246,500.

Determine (a) the break-even point in sales units and (b) the sales units required for the company to achieve a target profit of $54,230.

a. Break-even point in sales units ? units
b. Break-even point in sales units required for the company to achieve a target profit of $54,230 ? units
Answer #1

1.

A. Break even point = Fixed cost/Contribution margin per unit

= 516,375/(97-52)

= 11475 units

B. Break-even point = 516,375/(103 - 52)

= 10125 units

.

2

Breakeven point = 246,500/(150-65)

= 2900

Sales needed = (Fixed costs + Target profit)/Contribution margin per unit

= (246,500+54,230)/(150-65)

= 3538 units